A Pontiac Grand Am enters the Will Rogers Turnpike traveling at 50 miles per hour. One-half hour later, a Ford Mustang enters the turnpike at the same location and travels in the same direction at 80 miles per hour. How long will it take for the mustang to catch up to the Grand Am? When it catches up, how far will the two cars have traveled on the turnpike?
Please help! Thanks!

Answer by mananth(16949)

the distance traveled by both is same when one catches up with the other.
Let them meet at distance x
Grand Am speed= 50 mph
MUSTANG speed= 80 mph
hours
Time difference between the starting of the two= 0.50

Time taken by Grand Am = X / 50
Time taken by Car II = X / 80

X / 50 - X / 80 = 0.5

LCD - 400
Multiply equation by 400 400
400 / 50 x- x / 80 x=
8 x - 5 = 200
3 x = 200
/ 3
x= 66.67 miles